# MaskBit: Embedding-free Image Generation via Bit Tokens

### v1

@@ Abstract
Masked transformer models for class-conditional image generation have become a compelling alternative to diffusion models. Typically comprising two stages – an initial VQGAN model for transitioning between latent space and image space, and a subsequent Transformer model for image generation within latent space – these frameworks offer promising avenues for image synthesis. In this study, we present two primary contributions: Firstly, an empirical and systematic examination of VQGANs, leading to a modernized VQGAN. Secondly, a novel embedding-free generation network operating directly on bit tokens – a binary quantized representation of tokens with rich semantics. The first contribution furnishes a transparent, reproducible, and high-performing VQGAN model, enhancing accessibility and matching the performance of current state-of-the-art methods while revealing previously undisclosed details. The second contribution demonstrates that embedding-free image generation using bit tokens achieves a new state-of-the-art FID of 1.52 on the ImageNet $256\times 256$ benchmark, with a compact generator model of mere 305M parameters.
@@ 1 Motivation
Masked transformer models for class-conditioned image [4, 56] or text-to-image [5] generation have emerged as strong alternatives to auto-regressive models [11, 53, 29] and diffusion models [22, 43]. These masked transformer methods typically employ a two-stage framework: a discrete tokenizer (e.g., VQGAN [11]) projects the input from image space to a discrete, compact latent space, while a transformer model [50, 9] serves as the generator to create images in latent space from a masked token sequence. These methods achieve performance comparable to state-of-the-art auto-regressive and diffusion models, but with the advantage of requiring significantly fewer sampling steps and smaller model sizes, resulting in substantial speed-ups.
@@
Despite the success of masked transformer frameworks, the development details of a strong tokenizer have been largely overlooked. Moreover, one key aspect of training modern VQGAN-based tokenizers [4, 55, 56] – the perceptual loss – remains undiscussed. Since latent space-based generation relies on encoding and decoding the input image with the VQGAN tokenizer, the final image quality is influenced by both the generator network and the VQGAN. The importance of a publicly available, high-performance VQGAN model cannot be overstated, yet the most widely-used model and code-base still originate from the original work [11] developed over three years ago. Although stronger, closed-source VQGAN variants exist [4, 55, 56], their details are not fully shared in the literature, creating a significant performance gap for researchers without access to these advanced tokenizers. While the community has made attempts [3, 38] to reproduce these works, none have matched the performance (for both reconstruction and generation) reported in [4, 55, 56].
@@
In this paper, we undertake a systematic step-by-step study to elucidate the architectural design and training process necessary to create a modernized VQGAN model, referred to as VQGAN+. We provide a detailed ablation of key components in the VQGAN design, and propose several changes to them, including model and discriminator architecture, perceptual loss, and training recipe. As a result, we significantly enhance the VQGAN model, reducing the reconstruction FID from 7.94 [11] to 1.66, marking an impressive improvement of 6.28. Moreover, our modernized VQGAN+ improves the generation performance and establishes a strong, reproducible foundation for future generation frameworks based on VQGAN.
@@
Furthermore, we delve into embedding-free tokenization [56, 39], specifically implementing Lookup-Free Quantization (LFQ) [56] within our improved tokenizer framework. Our resulting method, employs a binary quantization process by projecting latent embeddings into $K$ dimensions and then quantizing them based on their sign values. This process produces bit tokens, where each token is represented by $K$ bits. We empirically observe that this representation captures high-level structured information, with bit tokens in close proximity being semantically similar. This insight leads us to propose a novel embedding-free generation model, MaskBit, which directly generates images using bit tokens, eliminating the need for learning new embeddings (from VQGAN token indices to new embedding values) as required in traditional VQGAN-based generators [11, 4, 56]. Consequently, MaskBit achieves state-of-the-art performance, with an FID score of 1.52 on the popular ImageNet $256\times 256$ image generation benchmark using a compact generator model of 305M parameters. All implementations will be made publicly available to support further research in this field.
@@ 2 Revisiting VQGAN
Image generation methods [11, 43] that operate in a latent space rather than in the pixel space require networks to project images into this latent space and then back to the pixel space. Such networks are usually trained with a reconstruction objective, as is common in VAE-inspired methods [27, 40, 11]. Over time, the training frameworks for these networks have become increasingly complex [11, 53, 4, 29, 55, 56]. Despite this complexity, the exploration and optimization of these frameworks remain significantly underrepresented in the scientific literature. In particular, issues in reproduction arise when crucial details are not transparently discussed, making a comparison under fair and consistent conditions hard. Moreover, fully disclosing all details enables the community to scrutinize and validate the true advancements brought by new designs. In response to this challenge, we conduct a systematic study on the essential modifications to the popular baseline Taming-VQGAN [11]. For this study, we use the term “VQGAN” to refer to any network of this type and “Taming-VQGAN” specifically to refer to the network published by [11]. In the following, we provide an empirical roadmap for developing a modernized VQGAN, named VQGAN+, aiming to bridge the performance gap and make advanced image generation techniques more accessible.
@@ 2.1 Preliminaries
Variational Autoencoders (VAEs) [27] compress high-dimensional data into low-dimensional representations using an encoder network and reverse this process with a decoder network. VAEs are typically trained with a reconstruction loss, such as the $L_{2}$ loss. However, minimizing the $L_{2}$ distance alone is insufficient for achieving visually satisfying results. Therefore, a perceptual loss term [25] based on the LPIPS metric [58] is used to reduce noise and improve image quality. Vector-Quantized VAEs (VQ-VAEs) [40] incorporate a learnable codebook that acts as a lookup table between the encoder and decoder. The codebook is trained with two $L_{2}$ losses: the commitment loss, which reduces the distance of the encoder output to the codebook entries, and the codebook loss, which minimizes the distance from the codebook entries to the encoder output. VQGANs [11] build on top of VQ-VAEs by incorporating a discriminator network to add an adversarial objective [15] to the training framework. The discriminator’s goal is to distinguish reconstructed images from original ones, while the VQ-VAE is trained to fool the discriminator.
@@ 2.2 The Experimental Setup
We follow standard practices to train and evaluate the network on ImageNet [8]. Unless specified otherwise, all networks are trained with a batch size of 256 for 300,000 iterations. Evaluation is performed using the reconstruction Fréchet Inception Distance (rFID) [18], where lower values indicate better performance. We use Taming-VQGAN [11] as our baseline and starting point, due to its thorough study and open-source availability, ensuring reproducibility of results. Taming-VQGAN is an encoder-decoder network that uses residual blocks [17] with convolutions for low-level feature processing and interleaved blocks of convolutions and self-attention [50] for high-level feature processing. In our experiments, we focus on encoder networks that produce latent embeddings with an output stride of 16, resulting in $16\times 16$ latent embeddings for input images of $256\times 256$ pixels. The learned vocabulary consists of 1,024 entries, each being a 256-dimensional vector.
@@ 2.3 VQGAN+: A Modern VQGAN
Figure 2: Detailed roadmap to build a modern VQGAN+. This overview summarizes the performance gains achieved by each proposed change to the architecture and training recipe. The reconstruction FID (rFID) is computed against the validation split of ImageNet at a resolution of 256. The popular and open-source Taming-VQGAN [11] serves as the baseline and starting point.
Starting with the baseline Taming-VQGAN [11] that attains 7.94 rFID on ImageNet-1K $256\times 256$ [8], we provide a step-by-step walk-through and ablation of all changes in the following paragraphs. A summary of the changes and their effects is given in Fig. 2.
Basic Training Recipe and Architecture.  The initial modifications to the Taming-VQGAN baseline are as follows: (1) removing attention blocks for a purely convolutional design, (2) adding symmetry to the generator and discriminator, and (3) updating the learning rate scheduler. Removing the attention layers, as adopted in recent methods [4, 55, 56], reduces computational complexity without sacrificing performance. We also observed discrepancies between the open-source implementation of Taming-VQGAN [11] and its description in the publication. The implementation uses two residual blocks per stage in the encoder but three per stage in the decoder. For our experiments, we adopt a symmetric architecture with two residual blocks per stage in both the encoder and decoder, as described in the publication. Additionally, we align the number of base channels in the generator and discriminator. Specifically, we reduce the channel dimension of the generator from 128 to 64 to match the discriminator. Here, base channels refer to the feature dimension in the first stage of the network. The third modification involves the learning rate scheduler. We replace a constant learning rate with a cosine-learning rate scheduler with warmup [36]. Specifically, we use a warmup phase of 5,000 iterations to reach the initial learning rate of $41e-$ 4 and then gradually decay it to 10% of the initial learning rate over time.
@@
Perceptual Loss.  The perceptual loss [25, 58] plays a crucial role in further reducing the rFID score. In Taming-VQGAN, the LPIPS [58] score obtained by the LPIPS VGG [46] network is minimized to improve image decoding. However, we noticed in the partially open-source code of recent work on image generation [55] [^1] https://github.com/google-research/magvit/, the usage of a ResNet50 [17] network instead to compute the perceptual loss. Although confirmed by the authors of [55, 56], this change has so far not been discussed in the literature. We believe this change was already introduced in earlier publications of the same group [4], but never transparently documented. Following the open-source code of [55], we apply an $L_{2}$ loss on the logits of a pretrained ResNet50 using the original and reconstructed images. While incorporating intermediate features into this loss can improve reconstruction further, it negatively affects generation performance in our experiments. Therefore, we compute the loss solely on the logits.
@@
Discriminator Update.  The original PatchGAN discriminator [11] employs $4\times 4$ convolutions and batch normalization [24], resulting in an output resolution of $30\times 30$ from a $256\times 256$ input and utilizing 11 million parameters (with 128 base channels). We propose to replace the $4\times 4$ convolution kernels with $3\times 3$ kernels and switch to group normalization [52], maintaining the same number of convolutions while producing a $32\times 32$ output resolution. We then apply $2\times 2$ max pooling to align the output stride between the generator and discriminator. These adjustments reduce the discriminator’s parameter count to 1.9 million, absorbing the earlier increase in parameters completely. In the second update, following prior work [56], we replace average pooling for downsampling with a precomputed $4\times 4$ Gaussian blur kernel using a stride of 2. Additionally, we incorporate LeCAM loss [49] to stabilize adversarial training.
@@
The Final Changes.  Our last addition to the training framework is the use of Exponential Moving Average (EMA) [42]. We found that EMA significantly stabilizes the training and improves convergence, while also providing a small performance boost. Orthogonal to the reconstruction performance, we also add an entropy loss [56] to ease the learning in the generation phase. To achieve the final performance of VQGAN+, we increase the number of training iterations from 300,000 to 1.35 million iterations, following common practices [56, 13].
@@
An Embedding-Free Variant.  Recent works [56, 39] have shown that removing the lookup table from the quantization [40] improves scalability. Following the Lookup-Free Quantization (LFQ) approach [56], we implement a binary quantization process by projecting the latent embeddings to $K$ dimensions ($K=12$ in this experiment) and then quantizing them based on their sign values. Due to this binary quantization, we can directly interpret the obtained quantized representations as their token numbers. We use the standard conversion between base-2 and base-10 integers (e.g., $1001_{2}=9_{10}$). In VQGAN+, the token is merely the index number of the embedding table; however, in this case, the token itself contains its corresponding representation, eliminating the need for an additional embedding. Implicitly, this defines a non-learnable codebook of size $2^{K}$. Given the compact nature of binary quantization with only $K$ feature dimensions, we do not anticipate significant gains in reconstruction. However, we will empirically demonstrate the clear advantages of this representation in the generation stage.
@@ 3 MaskBit: A New Embedding-free Image Generation Model
Since the typical pipeline of VQGAN-based image generation models [4, 56] includes two stages with Stage-I already discussed in the previous section, we focus on the Stage-II in this section. Stage-II refers to a transformer network that learns to generate images in the VQGAN’s latent space. During Stage-II training, the tokenizer produces latent representations of images, which are then masked according to a predefined schedule [4]. The masked tokens are fed into the Stage-II transformer network, which is trained with a categorical reconstruction objective for the masked tokens. This process resembles masked language modeling, making language models such as BERT [9] suitable for this stage. These transformer models “relearn” a new embedding for each token, as the input tokens themselves are just arbitrary index numbers.
Figure 3: Bit tokens exhibit structured semantic representations. We visualize a robustness test involving bit flipping. Specifically, we utilize our method to encode images into bit tokens, where each token is represented by $K$-bits ($K=12$ in this example). We then flip the $i$-th bit for all the bit tokens and reconstruct the images as usual. Interestingly, the reconstructed images from these bit-flipped tokens remain semantically consistent to the original image, exhibiting only minor visual modifications such as changes in texture, exposure, smoothness, color palette, or painterly quality.
Bit Tokens Are Semantically Structured.  Following our study of embedding-free tokenization, which eliminates the lookup table from Stage-I (i.e., removing the need for embedding tables in VQ-VAE [40]) by directly quantizing the tokens into $K$-bits, we study the achieved representation more carefully. As shown in [56], this representation yields high-fidelity image reconstruction results, indicating that images can be well represented and reconstructed using bit tokens. However, prior arts [56] have used different token spaces for Stage-I and Stage-II, intuitively allowing the Stage-II token to focus more on semantics, while enabling the Stage-I token space to capture all visual details necessary for faithful reconstruction. This raises the question:
@@
To address this, we first analyze the bit tokens’ learned representations for image reconstruction. Specifically, given a latent representation with 256 tokens from our model (where each token is represented by $K$-bits, and each bit by either $-1$ or $1$), we conduct a robustness test involving bit flipping. This test involves flipping the $i$-th bit (i.e., swapping $-1$ and $1$) for all 256 bit tokens and decoding them as usual to produce images. The bit-flipped tokens each have a Hamming distance of 1 from the original bit tokens. Surprisingly, we find that the reconstructed images from these bit-flipped tokens are still visually and semantically similar to the original images. This indicates that the representation of bit tokens has learned structured semantics, where neighboring tokens (within a Hamming distance of 1) are semantically similar to each other. We visualize an example of this in Fig. 3 and provide more examples in Sec. D. We note that this behaviour is not present in VQGAN methods. Conducting the same experiment with VQGAN+ leads to non-meaningful output, sharing no semantic or visual similarities with the original image. Since the generation model is usually semantically conditioned on external input such as classes, the model needs to learn the semantic relationships of tokens. Given this evidence of the inherent structured semantic representation in bit tokens, we believe that the learned rich semantic structure in the bit token representation can thus be a good representation for the Stage-II generation model.
@@
Figure 4: High-level overview of the architecture and comparison. Our training framework comprises two stages for image generation. In Stage-I, an encoder-decoder network compresses images into a latent representation and decodes them back. Stage-II masks the tokens, feeds them into a transformer and predicts the masked tokens. Most prior art uses VQGAN-based methods (top) that learn independent embedding tables in both stages. In VQGAN-based methods, only indices of embedding tables are shared across stages, but not the embeddings. In MaskBit, however, neither Stage-I nor Stage-II utilizes embedding tables. The Stage-I predicts bit tokens by using binary quantization on the encoder output directly. The Stage-II partitions the shared bit tokens into groups (e.g., 2 groups), masks and feeds them into a transformer, and predicts the masked bit tokens.
@@
Proposing MaskBit.  This motivates us to propose MaskBit for Stage-II, a transformer-based generative model that generates images directly using bit tokens, removing the need for the embedding lookup table as required by existing methods [4, 56]. Our training framework is illustrated in Fig. 4. Prior art using VQGAN approaches share only indices of the embedding tables between the two stages, but learn independent embeddings. Taking advantage of the built-in semantic structure of bit tokens, the proposed MaskBit can share the tokens directly between Stage-I and Stage-II. We discuss more detailed differences between the two approaches in Sec. C. In the next paragraph, we provide details on bit token masking.
@@
Masked Bits Modeling.  The Stage-II training follows the masked modeling framework [9], where a certain number of tokens are masked (i.e., replaced with a special mask token) before being fed into the transformer, which is trained to recover the masked tokens. This approach requires an additional entry in the embedding table to learn the embedding vector for the special mask token. However, this presents a challenge for an embedding-free setup, where images are generated directly using bit tokens without embedding lookup. Specifically, it raises the question of how to represent the masked bit tokens in the new framework. To address this challenge, we propose a straightforward yet effective solution: using zeros to represent the masked bit tokens. In particular, a bit token $t$ is represented as $t\in\{-1,1\}^{K}$ (i.e., $K$-bits, with each bit being either $-1$ or $1$), while we set all masked bits to zero. Consequently, these masked bit tokens do not contribute to the image representation.

### v2

@@ Abstract
Masked transformer models for class-conditional image generation have become a compelling alternative to diffusion models. Typically comprising two stages – an initial VQGAN model for transitioning between latent space and image space, and a subsequent Transformer model for image generation within latent space – these frameworks offer promising avenues for image synthesis. In this study, we present two primary contributions: Firstly, an empirical and systematic examination of VQGANs, leading to a modernized VQGAN. Secondly, a novel embedding-free generation network operating directly on bit tokens – a binary quantized representation of tokens with rich semantics. The first contribution furnishes a transparent, reproducible, and high-performing VQGAN model, enhancing accessibility and matching the performance of current state-of-the-art methods while revealing previously undisclosed details. The second contribution demonstrates that embedding-free image generation using bit tokens achieves a new state-of-the-art FID of 1.52 on the ImageNet $256\times 256$ benchmark, with a compact generator model of mere 305M parameters. The code for this project is available on https://github.com/markweberdev/maskbit.
@@ 1 Motivation
Masked transformer models for class-conditioned image (Chang et al., 2022; Yu et al., 2024a) or text-to-image (Chang et al., 2023) generation have emerged as strong alternatives to auto-regressive models (Esser et al., 2021; Yu et al., 2022a; Lee et al., 2022a) and diffusion models (Ho et al., 2020; Rombach et al., 2022). These masked transformer methods typically employ a two-stage framework: a discrete tokenizer (e.g., VQGAN (Esser et al., 2021)) projects the input from image space to a discrete, compact latent space, while a transformer model (Vaswani et al., 2017; Devlin et al., 2018) serves as the generator to create images in latent space from a masked token sequence. These methods achieve performance comparable to state-of-the-art auto-regressive and diffusion models, but with the advantage of requiring significantly fewer sampling steps and smaller model sizes, resulting in substantial speed-ups.
@@
Despite the success of masked transformer frameworks, the development details of a strong tokenizer have been largely overlooked. Moreover, one key aspect of training modern VQGAN-based tokenizers (Chang et al., 2022; Yu et al., 2023; 2024a) – the perceptual loss – remains undiscussed. Since latent space-based generation relies on encoding and decoding the input image with the VQGAN tokenizer, the final image quality is influenced by both the generator network and the VQGAN. The importance of a publicly available, high-performance VQGAN model cannot be overstated, yet the most widely-used model and code-base still originate from the original work (Esser et al., 2021) developed over three years ago. Although stronger, closed-source VQGAN variants exist (Chang et al., 2022; Yu et al., 2023; 2024a), their details are not fully shared in the literature, creating a significant performance gap for researchers without access to these advanced tokenizers. While the community has made attempts (Besnier & Chen, 2023; Luo et al., 2024) to reproduce these works, none have matched the performance (for both reconstruction and generation) reported in (Chang et al., 2022; Yu et al., 2023; 2024a).
@@
In this paper, we undertake a systematic step-by-step study to elucidate the architectural design and training process necessary to create a modernized VQGAN model, referred to as VQGAN+. We provide a detailed ablation of key components in the VQGAN design, and propose several changes to them, including model and discriminator architecture, perceptual loss, and training recipe. As a result, we significantly enhance the VQGAN model, reducing the reconstruction FID from 7.94 (Esser et al., 2021) to 1.66, marking an impressive improvement of 6.28. Moreover, our modernized VQGAN+ improves the generation performance and establishes a strong, reproducible foundation for future generation frameworks based on VQGAN.
@@
Furthermore, we delve into embedding-free tokenization (Yu et al., 2024a; Mentzer et al., 2024), specifically implementing Lookup-Free Quantization (LFQ) (Yu et al., 2024a) within our improved tokenizer framework. Our resulting method, employs a binary quantization process by projecting latent embeddings into $K$ dimensions and then quantizing them based on their sign values. This process produces bit tokens, where each token is represented by $K$ bits. We empirically observe that this representation captures high-level structured information, with bit tokens in close proximity being semantically similar. This insight leads us to propose a novel embedding-free generation model, MaskBit, which directly generates images using bit tokens, eliminating the need for learning new embeddings (from VQGAN token indices to new embedding values) as required in traditional VQGAN-based generators (Esser et al., 2021; Chang et al., 2022; Yu et al., 2024a). Consequently, MaskBit achieves state-of-the-art performance, with an FID score of 1.52 on the popular ImageNet $256\times 256$ image generation benchmark using a compact generator model of 305M parameters. All implementations will be made publicly available to support further research in this field.
@@ 2 Revisiting VQGAN
Image generation methods (Esser et al., 2021; Rombach et al., 2022) that operate in a latent space rather than in the pixel space require networks to project images into this latent space and then back to the pixel space. Such networks are usually trained with a reconstruction objective, as is common in VAE-inspired methods (Kingma & Welling, 2014; Oord et al., 2017; Esser et al., 2021). Over time, the training frameworks for these networks have become increasingly complex (Esser et al., 2021; Yu et al., 2022a; Chang et al., 2022; Lee et al., 2022a; Yu et al., 2023; 2024a). Despite this complexity, the exploration and optimization of these frameworks remain significantly underrepresented in the scientific literature. In particular, issues in reproduction arise when crucial details are not transparently discussed, making a comparison under fair and consistent conditions hard. Moreover, fully disclosing all details enables the community to scrutinize and validate the true advancements brought by new designs. In response to this challenge, we conduct a systematic study on the essential modifications to the popular baseline Taming-VQGAN (Esser et al., 2021). For this study, we use the term “VQGAN” to refer to any network of this type and “Taming-VQGAN” specifically to refer to the network published by (Esser et al., 2021). In the following, we provide an empirical roadmap for developing a modernized VQGAN, named VQGAN+, aiming to bridge the performance gap and make advanced image generation techniques more accessible.
@@ 2.1 Preliminaries
Variational Autoencoders (VAEs) (Kingma & Welling, 2014) compress high-dimensional data into low-dimensional representations using an encoder network and reverse this process with a decoder network. VAEs are typically trained with a reconstruction loss, such as the $L_{2}$ loss. However, minimizing the $L_{2}$ distance alone is insufficient for achieving visually satisfying results. Therefore, a perceptual loss term (Johnson et al., 2016) based on the LPIPS metric (Zhang et al., 2018) is used to reduce noise and improve image quality. Vector-Quantized VAEs (VQ-VAEs) (Oord et al., 2017) incorporate a learnable codebook that acts as a lookup table between the encoder and decoder. The codebook is trained with two $L_{2}$ losses: the commitment loss, which reduces the distance of the encoder output to the codebook entries, and the codebook loss, which minimizes the distance from the codebook entries to the encoder output. VQGANs (Esser et al., 2021) build on top of VQ-VAEs by incorporating a discriminator network to add an adversarial objective to the training framework (Goodfellow et al., 2014). The discriminator’s goal is to distinguish reconstructed images from original ones, while the VQ-VAE is trained to fool the discriminator.
@@ 2.2 The Experimental Setup
We follow standard practices to train and evaluate the network on ImageNet (Deng et al., 2009). Unless specified otherwise, all networks are trained with a batch size of 256 for 300,000 iterations. Evaluation is performed using the reconstruction Fréchet Inception Distance (rFID) (Heusel et al., 2017), where lower values indicate better performance. We use Taming-VQGAN (Esser et al., 2021) as our baseline and starting point, due to its thorough study and open-source availability, ensuring reproducibility of results. Taming-VQGAN is an encoder-decoder network that uses residual blocks (He et al., 2016) with convolutions for low-level feature processing and interleaved blocks of convolutions and self-attention (Vaswani et al., 2017) for high-level feature processing. In our experiments, we focus on encoder networks that produce latent embeddings with an output stride of 16, resulting in $16\times 16$ latent embeddings for input images of $256\times 256$ pixels. The learned vocabulary consists of 1,024 entries, each being a 256-dimensional vector.
@@ 2.3 VQGAN+: A Modern VQGAN
Figure 2: Detailed roadmap to build a modern VQGAN+. This overview summarizes the performance gains achieved by each proposed change to the architecture and training recipe. The reconstruction FID (rFID) is computed against the validation split of ImageNet at a resolution of 256. The popular and open-source Taming-VQGAN (Esser et al., 2021) serves as the baseline and starting point.
Starting with the baseline Taming-VQGAN (Esser et al., 2021) that attains 7.94 rFID on ImageNet-1K $256\times 256$ (Deng et al., 2009), we provide a step-by-step walk-through and ablation of all changes in the following paragraphs. A summary of the changes and their effects is given in Fig. 2.
Basic Training Recipe and Architecture.  The initial modifications to the Taming-VQGAN baseline are as follows: (1) removing attention blocks for a purely convolutional design, (2) adding symmetry to the generator and discriminator, and (3) updating the learning rate scheduler. Removing the attention layers, as adopted in recent methods (Chang et al., 2022; Yu et al., 2023; 2024a), reduces computational complexity without sacrificing performance. We also observed discrepancies between the open-source implementation of Taming-VQGAN (Esser et al., 2021) and its description in the publication. The implementation uses two residual blocks per stage in the encoder but three per stage in the decoder. For our experiments, we adopt a symmetric architecture with two residual blocks per stage in both the encoder and decoder, as described in the publication. Additionally, we align the number of base channels in the generator and discriminator. Specifically, we reduce the channel dimension of the generator from 128 to 64 to match the discriminator. Here, base channels refer to the feature dimension in the first stage of the network. The third modification involves the learning rate scheduler. We replace a constant learning rate with a cosine-learning rate scheduler with warmup (Loshchilov & Hutter, 2016). Specifically, we use a warmup phase of 5,000 iterations to reach the initial learning rate of $41e-$ 4 and then gradually decay it to 10% of the initial learning rate over time.
@@
Perceptual Loss.  The perceptual loss (Johnson et al., 2016; Zhang et al., 2018) plays a crucial role in further reducing the rFID score. In Taming-VQGAN, the LPIPS (Zhang et al., 2018) score obtained by the LPIPS VGG (Simonyan & Zisserman, 2015) network is minimized to improve image decoding. However, we noticed in the partially open-source code of recent work on image generation (Yu et al., 2023) [^1] https://github.com/google-research/magvit/, the usage of a ResNet50 (He et al., 2016) network instead to compute the perceptual loss. Although confirmed by the authors of (Yu et al., 2023; 2024a), this change has so far not been discussed in the literature. We believe this change was already introduced in earlier publications of the same group (Chang et al., 2022), but never transparently documented. Following the open-source code of (Yu et al., 2023), we apply an $L_{2}$ loss on the logits of a pretrained ResNet50 using the original and reconstructed images. While incorporating intermediate features into this loss can improve reconstruction further, it negatively affects generation performance in our experiments. Therefore, we compute the loss solely on the logits.
@@
Discriminator Update.  The original PatchGAN discriminator (Esser et al., 2021) employs $4\times 4$ convolutions and batch normalization (Ioffe & Szegedy, 2015), resulting in an output resolution of $30\times 30$ from a $256\times 256$ input and utilizing 11 million parameters (with 128 base channels). We propose to replace the $4\times 4$ convolution kernels with $3\times 3$ kernels and switch to group normalization (Wu & He, 2018), producing a $16\times 16$ output resolution to align the output stride between the generator and discriminator. These adjustments reduce the discriminator’s parameter count to 7.4 million. In the second update, following prior work (Yu et al., 2024a), we replace average pooling for downsampling with a precomputed $4\times 4$ Gaussian blur kernel using a stride of 2 (Zhang, 2019). Additionally, we incorporate LeCAM loss (Tseng et al., 2021) to stabilize adversarial training.
@@
The Final Changes.  Our last addition to the training framework is the use of Exponential Moving Average (EMA) (Polyak & Juditsky, 1992). We found that EMA significantly stabilizes the training and improves convergence, while also providing a small performance boost. Orthogonal to the reconstruction performance, we also add an entropy loss (Yu et al., 2024a) to ease the learning in the generation phase. To achieve the final performance of VQGAN+, we increase the number of training iterations from 300,000 to 1.35 million iterations, following common practices (Yu et al., 2024a; Gao et al., 2023).
@@
An Embedding-Free Variant.  Recent works (Yu et al., 2024a; Mentzer et al., 2024) have shown that removing the lookup table from the quantization Oord et al. (2017) improves scalability. Following the Lookup-Free Quantization (LFQ) approach (Yu et al., 2024a), we implement a binary quantization process by projecting the latent embeddings to $K$ dimensions ($K=12$ in this experiment) and then quantizing them based on their sign values. Due to this binary quantization, we can directly interpret the obtained quantized representations as their token numbers. We use the standard conversion between base-2 and base-10 integers (e.g., $1001_{2}=9_{10}$). In VQGAN+, the token is merely the index number of the embedding table; however, in this case, the token itself contains its corresponding representation, eliminating the need for an additional embedding. Implicitly, this defines a non-learnable codebook of size $2^{K}$. Given the compact nature of binary quantization with only $K$ feature dimensions, we do not anticipate significant gains in reconstruction. However, we will empirically demonstrate the clear advantages of this representation in the generation stage.
@@ 3 MaskBit: A New Embedding-free Image Generation Model
Since the typical pipeline of VQGAN-based image generation models (Chang et al., 2022; Yu et al., 2024a) includes two stages with Stage-I already discussed in the previous section, we focus on the Stage-II in this section. Stage-II refers to a transformer network that learns to generate images in the VQGAN’s latent space. During Stage-II training, the tokenizer produces latent representations of images, which are then masked according to a predefined schedule (Chang et al., 2022). The masked tokens are fed into the Stage-II transformer network, which is trained with a categorical reconstruction objective for the masked tokens. This process resembles masked language modeling, making language models such as BERT (Devlin et al., 2018) suitable for this stage. These transformer models “relearn” a new embedding for each token, as the input tokens themselves are just arbitrary index numbers.
Figure 3: Bit tokens exhibit structured semantic representations. We visualize images obtained after corrupting tokens via bit flipping. Specifically, we utilize our method to encode images into bit tokens, where each token is represented by $K$-bits ($K=12$ in this example). We then flip the $i$-th bit for all the bit tokens and reconstruct the images as usual. Interestingly, the reconstructed images from these bit-flipped tokens remain semantically consistent to the original image, exhibiting only minor visual modifications such as changes in texture, exposure, smoothness, color palette, or painterly quality.
Bit Tokens Are Semantically Structured.  Following our study of embedding-free tokenization, which eliminates the lookup table from Stage-I (i.e., removing the need for embedding tables in VQ-VAE (Oord et al., 2017)) by directly quantizing the tokens into $K$-bits, we study the achieved representation more carefully. As shown in (Yu et al., 2024a), this representation yields high-fidelity image reconstruction results, indicating that images can be well represented and reconstructed using bit tokens. However, prior arts (Yu et al., 2024a) have used different token spaces for Stage-I and Stage-II, intuitively allowing the Stage-II token to focus more on semantics, while enabling the Stage-I token space to capture all visual details necessary for faithful reconstruction. This raises the question:
@@
To address this, we first analyze the bit tokens’ learned representations for image reconstruction. Specifically, given a latent representation with 256 tokens from our model (where each token is represented by $K$-bits, and each bit by either $-1$ or $1$), we conduct a a bit flipping experiment to show how the structure of the token space relates to the content. This test involves flipping the $i$-th bit (i.e., swapping $-1$ and $1$) for all 256 bit tokens and decoding them as usual to produce images. The bit-flipped tokens each have a Hamming distance of 1 from the original bit tokens. Surprisingly, we find that the reconstructed images from these bit-flipped tokens are still visually and semantically similar to the original images. This indicates that the representation of bit tokens has learned structured semantics, where neighboring tokens (within a Hamming distance of 1) are semantically similar to each other. We visualize an example of this in Fig. 3 and provide more examples in Sec. E. We note that the structural differences in the latent space between bit tokens and learnable embeddings also have implications on the generator’s representation, which we describe in Appendix D. Since the generation model is usually semantically conditioned on external input such as classes, the model needs to learn the semantic relationships of tokens. Given this evidence of the inherent structured semantic representation in bit tokens, we believe that the learned rich semantic structure in the bit token representation can thus be a good representation for the Stage-II generation model.
@@
Figure 4: High-level comparison of the architectures. Our training framework comprises two stages for image generation. In Stage-I, an encoder-decoder network compresses images into a latent representation and decodes them back. Stage-II masks the tokens, feeds them into a transformer and predicts the masked tokens. Most prior art uses VQGAN-based methods (top) that learn independent embedding tables in both stages. In VQGAN-based methods, only indices of embedding tables are shared across stages, but not the embeddings. In MaskBit, however, neither Stage-I nor Stage-II utilizes embedding tables. The Stage-I predicts bit tokens by using binary quantization on the encoder output directly. The Stage-II partitions the shared bit tokens into groups, masks and feeds them into a transformer to predict the masked bit tokens. An illustration of grouping and masking of bit tokens can be found in the Appendix C.
@@
Proposing MaskBit.  This motivates us to propose MaskBit for Stage-II, a transformer-based generative model that generates images directly using bit tokens, removing the need for the embedding lookup table as required by existing methods (Chang et al., 2022; Yu et al., 2024a). Our training framework is illustrated in Fig. 4. Prior art using VQGAN approaches share only indices of the embedding tables between the two stages, but learn independent embeddings. Taking advantage of the built-in semantic structure of bit tokens, the proposed MaskBit can share the tokens directly between Stage-I and Stage-II. We discuss more detailed differences between the two approaches in Sec. D. Next, we provide details on bit token masking.
@@
Masked Bits Modeling.  The Stage-II training follows the masked modeling framework (Devlin et al., 2018), where a certain number of tokens are masked (i.e., replaced with a special mask token) before being fed into the transformer, which is trained to recover the masked tokens. This approach requires an additional entry in the embedding table to learn the embedding vector for the special mask token. However, this presents a challenge for an embedding-free setup, where images are generated directly using bit tokens without embedding lookup. Specifically, it raises the question of how to represent the masked bit tokens in the new framework. To address this challenge, we propose a straightforward yet effective solution: using zeros to represent the masked bit tokens. In particular, a bit token $t$ is represented as $t\in\{-1,1\}^{K}$ (i.e., $K$-bits, with each bit being either $-1$ or $1$), while we set all masked bits to zero. Consequently, these masked bit tokens do not contribute to the image representation.